a particle moving with a initial velocity u and passes succesfully over 2 halves of a given distance with acceleraton f and f1 respectively.Prove that the final velocity is equal to the initial velocity if the entire distance was traversed with a uniform accleration of (f+f1)/2

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
Let the half distance = x
For the first part travel
v2=u2+2xfx/2
For the second part travel,
vf2=u2+2fx+2f1x
 
=>vf=√(u^2+2fx+2f1x)
 
when avg acceleration is (f+f1)/2
 
vf2=u2+2(f+f1)x2x/2
=> vf=√(u^2+2fx+2f1x)
 
Hence equal. Proved.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)